### v2

# CamC2V: Context-aware Controllable Video Generation

Abstract
Recently, image-to-video (I2V) diffusion models have demonstrated impressive scene understanding and generative quality, incorporating image conditions to guide generation. However, these models primarily animate static images without extending beyond their provided context. Introducing additional constraints, such as camera trajectories, can enhance diversity but often degrade visual quality, limiting their applicability for tasks requiring faithful scene representation. We propose CamC2V, a context-to-video (C2V) model that integrates multiple image conditions as context with 3D constraints alongside camera control to enrich both global semantics and fine-grained visual details. This enables more coherent and context-aware video generation. Moreover, we motivate the necessity of temporal awareness for an effective context representation. Our comprehensive study on the RealEstate10K dataset demonstrates improvements in visual quality and camera controllability. We will publish our code upon acceptance.
1 Introduction
Diffusion models have become a prominent approach for video generation producing high-quality videos based on user inputs. To make such approaches attractive for digital content creation, controllability achieved through specific conditioning of the generations, like human poses [27, 19], style [10, 30], motion [23, 20] or camera trajectories [33, 6, 26] have been a widely studied topic.
While text-to-video (T2V) diffusion models like VideoCrafter [2] or CogVideoX [29] have full freedom over the visual design, more recent image-to-video (I2V) models employ an image to convey style and scene context. Due to the typically short duration ( seconds) of the generated videos, the image provides sufficient context to define the scene to render. With the ultimate objective of matching the generative quality and capabilities of traditional rendering engines, these approaches still require further development to achieve a fine-grained control over style, motion and scene composition, to allow for fully customizable video creation.
As illustrated in Fig. 1, the initial reference frame alone provides only limited context for the diffusion process. Once the camera pans, the visual quality degrades and arbitrary interpretations of the scene by the diffusion model become evident. To address this, we introduce CamC2V, a novel conditioning mechanism that allows users to supply multiple context views, ensuring a comprehensive definition of the scene in which the video is generated.
Our proposed Context-aware Encoder integrates these context views into two complementary streams: a high-level semantic stream and a 3D-aware visual stream. This dual-stream approach provides the diffusion model with both a global semantic context and a detailed pixel-level visual embedding. By inserting 3D geometric constraints in the feature aggregation, we effectively retrieve important features from the context while filtering out irrelevant ones. This allows our method to considerably enhance the visual coherence of existing approaches. In summary, our key contributions are as follows.
We propose CamC2V, a camera-controllable context-aware diffusion model, which conditions the diffusion process on multiple context frames through a dual-stream encoder retrieving high-level semantic features and low-level visual cues from the context.
We introduce a 3D-aware cross-attention mechanism leveraging epipolar constraints to effectively retrieve context from posed images.
Our temporally-aware embedding strategy better aligns the context at different frame timesteps.
Our method achieves a 24.09% improvement in visual quality over the state-of-the-art methods on the RealEstate10K dataset.
2 Related Works
Diffusion-based Video Generation.
Originally developed for image generation [7, 15], diffusion models have since demonstrated great success synthesizing high-quality videos [8, 1]. Models such as SVD [1], LAVIE [21] or VideoCrafter [2] have shown great success in distilling text-to-video (T2V) diffusion models from text-to-image (T2I) diffusion models by inserting temporal attention blocks modeling the added time dimension. Building on top, models like DynamiCrafter [25], Seine [3] or I2vgen-XL [32] further fine-tune these models for image-to-video (I2V) generation showing impressive results.
Camera-controllable Video Generation.
Concurrent work also focuses on adding camera control to diffusion models allowing the user to define the trajectory along a video is generated. While initial work such as MotionCtrl [23], AnimateDiff [5] or Direct-a-Video [28] model camera movements through camera-motion primitives, recent approaches such as CameraCtrl [6], CamCo [26] or CamI2V [33] directly insert the camera poses showcasing fine-grained camera control. A key is the dense supervisory signal, such as pixel-wise camera rays represented through Plücker coordinates, which are encoded and inserted into the diffusion model in a ControlNet-like fashion [31].
CamCo and CamI2V further demonstrate that epipolar geometry can serve as an effective constraint in the information aggregation of vanilla attention mechanism. While CamCo employs cross-attention to constraint the feature aggregation from the condition frame, CamI2V constrain the temporal self-attention itself to guide the diffusion process and thus improving the 3D consistency and camera trajectory.
Multi-Image Condition.
Large camera movements or longer generations result in multiple scenes being generated in one video which is insufficiently represented through a singular reference image typically employed in concurrent image-to-video diffusion models [25, 3, 32, 29]. Models like Gen-L-Video [18], MEVG [13] or VideoStudio [11] explore the insertion of multiple text prompts to give a broader context across the temporal domain for longer video generation. This is achieved by generating distinct short videos with different text conditions and optimizing the the noise between them either in a divide-and-conquer or auto-regressive setup to generate long consistent videos. Other approaches like 4DiM [24] or Seine [3] explore the insertion of multiple image conditions to induce motion cues or as key-frames to interpolate between.
3 Preliminaries
Before we describe in Section 4 our novel method, which enhances the context-awareness of pre-trained diffusion models by conditioning on multiple context views rather than a single reference frame, we briefly describe components of our baseline model, CamI2V, which extends DynamiCrafter [25], a latent image-to-video diffusion model with camera pose conditioning.
Latent Video Diffusion Models.
Latent video diffusion models learn a latent video data distribution by gradually reconstructing noisy latents sampled from a Gaussian distribution:
where hyperparameters determine the level of noise added at each timestep. The latent space is defined through a pre-trained auto-encoder, e.g. a pre-trained VQGAN [4] for DynamiCrafter, consisting of an encoder and a decoder . Conditioned on a text condition and a reference image , the diffusion model is then trained to predict the noise at timestep using a simple reconstruction loss:
The diffusion model itself is typically implemented as a UNet, e.g. a 3D-Unet [35] in DynamiCrafter, where denotes the neural network’s parameters.
Camera Conditioning.
To incorporate camera control, CamI2V employs a dense supervisory signal using pixel-wise embeddings of camera rays, represented via Plücker coordinates. Specifically, for each pixel the Plücker coordinates are computed using the normalized ray direction and the ray origin (the camera focal point).
The ray direction relative to a reference coordinate frame—such as the camera coordinate system of the initial frame—is derived from the intrinsics and extrinsics as:
These embeddings are further encoded at multiple resolutions and integrated into the epipolar attention blocks inserted into the U-Net.
4 Method
Image-to-video diffusion models generate videos based on a single reference frame and an optional text condition . Additionally, camera-controlled diffusion models are conditioned on a camera trajectory allowing precise control of the camera view at each timestep. The reference frame does not always provide the necessary context corresponding to the camera trajectory. This can lead to insufficient visual quality of the generated frames. In contrast, we propose a new scheme coined context-to-video which enhances the generation process with a rich context conveyed through additional context frames and their poses
Our Context-aware Encoder, shown in Fig. 2, extends DynamiCrafter’s Dual-stream Image Injection to support multiple image conditions. Natively, it conditions the model at the pixel level by concatenating reference latents with noisy latents along the channel dimension, which restricts the generations to the narrow context provided by the reference image. Additionally, to better guide the diffusion process, semantic features aggregated from CLIP-embedded image and text conditions are integrated layer-wise through spatial cross-attention. To utilize the pre-trained generative capabilities of the diffusion model and refrain from fine-tuning large parts of the U-Net, we chose to inject our condition in those streams.
Semantic Stream.
We adopt DynamiCrafter’s query transformer to integrate cross-modal information from the CLIP-embedded reference image , the text condition , and additional context frames . Specifically, employs learnable latent query tokens to gather context across multiple layers of cross-attention and feed-forward networks, yielding a global representation:
To preserve strong cross-modal context aggregation, we initialize from DynamiCrafter’s Dual-stream Image Injection module and fine-tune it to handle multiple image conditions.
Visual Stream.
While the semantic stream provides a well-suited global context representation, it lacks fine-grained visual details due to CLIP’s inherent training on visual-language alignment, which favors high-level representations of single entities.
To enhance context-aware generation, we integrate our visual condition directly into DynamiCrafter’s image conditioning. Specifically, we embed the context frames into the latent space and introduce pixel-wise learnable context tokens . The context tokens serve as queries in a query transformer, similar to the semantic stream, to aggregate timestep- and pixel-wise features from the latent context frames.
3D Awareness.
To introduce 3D awareness, we employ an epipolar cross-attention mechanism which guides the feature aggregation to only consider potentially relevant features. Specifically, each token , illustrated in Fig. 3, describes a pixel at timestep . Employing the provided camera pose at the given timestep, we can compute the epipolar line in each context view . Using the point-to-line distance:
we produce the epipolar mask masking out pixels with a distance larger than a threshold , set to half of the diagonal of the latent feature space, in the cross-attention mechanism:
where describes the learnable context queries and the latent embedded context frames.
Temporal Awareness.
The native pixel-level embedding of DynamiCrafter is agnostic to the timestep within the video as each timestep is provided with the same condition. Thus, to further enforce the diffusion model to attend to context provided at specific timesteps, we found it advantageous to employ a sinusoidal timestep embedding. In practice, we concatenate the timestep embedding to our context embeddings before forwarding it through a feed-forward network.
Finally, the visual stream of our Context-aware Encoder maps a spatially distributed embedding represented through the latent embedding of posed views to a timestep-wise embedding:
To retain the reference image as a strong anchor to the generation and smoothly insert the new condition, we employ a 3D zero-convolution which weighs the usage of DynamiCrafter’s native condition and ours before adding them together.
Log-weighted Loss.
Our Context-aware Encoder injects crucial information, especially to later frames, that have limited context in the baseline methods leading to degrading visual quality. To force the training process to focus on such frames that rely most heavily on our context‐aware conditioning, we apply a logarithmic re‐weighting to the standard reconstruction loss along the time axis. Specifically, for each frame in the sequence, we define:
This not only improves generative quality but also stabilizes the training, mitigating divergence in later stages of training.
5 Experiments
| Method | FVD | MSE | TransErr | RotErr | CamMC | |
| VideoGPT | StyleGAN | |||||
| MotionCtrl | 78.30 | 64.47 | 3654.54 | 2.89 | 2.04 | 4.34 |
[…]
| CamI2V | 71.01 | 57.90 | 2692.84 | 1.79 | 1.16 | 2.58 |
[…]
In the following, we thoroughly analyze our method in the setup described in Sec. 5.1. Sec. 5.2 and Sec. 5.3 compares our method to baseline camera-controlled methods and argue the effectiveness of the provided context. Next, in Sec. 5.4, we investigate our design choices and motivate the importance of 3D and temporal awareness in our method and the complementary effect of our two-stream design. Finally, in Sec. 5.5 we compare our method to a state-of-the-art Novel View Synthesis approach, employing similar conditioning signals.
5.1 Setup
Dataset
The RealEstate10K [34] comprises approximately 70K video clips at 720p of static scenes depicting indoor and outdoor house tours. The clips are annotated with camera extrinsic and intrinsic values obtained through the ORB-SLAM2 [12] pipeline. Additionally, we use the captions provided by the authors of CameraCtrl [6]. The video clips are then center-cropped to a size of and clipped to short frames of length with a stride sampled between 1 and 10.
Metrics
We evaluate our method with respect to generative quality, the faithfullness to the provided context and the camera trajectory. Firstly, to ensure improved visual quality we report the Frechet Video Distance (FVD) [17]. To ensure the faithfulness with respect to the additional context, we evaluate the pixel-wise mean squared error (MSE) and the Structural Similarity Index (SSIM) [22] independently for each timestep.
Finally, to examine the generated camera trajectory we follow the evaluation paradigm proposed by CameraCtrl and CamI2V. Using the structure-from-motion pipeline GLOMAP [14], we estimated the camera rotation and translation for each camera and compute the independent rotation and translation errors, RotErr and TransErr respectively, as well as the combined element-wise error CamMC:
| RotErr | (9) | |||
| TransErr | (10) | |||
| CamMC | (11) |
To counteract the randomness of GLOMAP, we follow the scheme to average the metrics over successful runs out of 5 trials. All metrics are computed on a subset consisting of videos extending over a duration of over seconds to ensure sufficient additional context to be sampled from and prevent avoid sampling to close to the frame clip.
Implementation Details
Our model builds on DynamiCrafter with inserted camera control from CamI2V. Training is initialized from CamI2V checkpoints at 50K iterations, freezing all parameters except for our Context-aware Encoder. We train for 50K iterations at a resolution of , using the Adam optimizer with a fixed learning rate of and a batch size of . Following the baseline, we chose Lightning as our training framework with mixed-precision using DeepSpeed ZeRo-1 on 4 NVIDIA A100 GPUs for approximately 7 days. For comparison to the baseline methods we use the re-implementations of MotionCtrl [23] and CameraCtrl [6] provided by the authors of CamI2V. We sample 1-4 context frames uniformly from the complete videos during training. For fairness, during evaluation we only sample from outside the frames window to be generated.
5.2 Quantitative Comparison
To show the effectiveness of the additional context provided by our method, we compare against several camera-controlled methods, namely MotionCtrl [23], CameraCtrl [6] and CamI2V [33]. Tab. 1 presents the comparison of our method against the baseline methods. Our model achieves an improvement of in terms of the FVD score highlighting the effectiveness of added context for video generation.
To further evaluate the context-awareness of our method, we report the MSE in Fig. 4 between the generated videos and the ground-truth videos on a per-frame basis to assess the improvement especially for later frames that typically lack sufficient context from the reference frame. Additionally, to assess the visual quality of each frame, we compute the SSIM metric on each frame.
It is visible that the visual quality degrades logarithmically with the video length as the diffusion model lacks sufficient context. Our method outperforms the baseline methods in both MSE and SSIM, especially for later frames. This shows that providing the diffusion process with additional context can stabilize the generative quality over time.
Additionally, we investigate the accuracy of the generated camera trajectory with respect to the RotErr, TransErr and CamMC. We observe a slightly improved rotational error compared to CamI2V’s, indicating an improved camera trajectory of our method. As the evaluation pipeline, GLOMAP, used for estimating the camera trajectory matches keypoint features to simultaneously estimate the camera trajectory and reconstruct a 3D scene using bundle adjustment and we do not train the camera encoder, nor the diffusion model itself, this improved camera trajectory is mainly linked to an improved 3D consistency and visual quality of the generated scene. This demonstrates that the additionally provided context enforces are more faithful representation of the 3D scene.
5.3 Qualitative Comparison
Fig. 5 shows different samples from our method compared against CamI2V. It is evident that the reference frame does not provide sufficient context for the generation past the first few frames. This results in visually degrading image quality and unrealistic generations of the baseline method.
In contrast, our method is provided with an additional context frame sampled from a later timestep past the window frame that shows entities outside of the field of view of the reference frame or obstructed by obstacles. Our method is able to comprehend the position of these entities in space and effectively embed it into the timestep-wise embedding resulting in these objects being placed at correct locations in later frames. Moreover, it is visible, while the baseline method produces artifacts not visible in its condition, the extended context provides an additional constraint preventing unwanted artifacts.
5.4 Ablation Studies
| Multi-Cond. | Epipolar | Time | FVD | MSE | ||||
| Pixel | Sem. | VideoGPT | StyleGAN | Total | t=2 | t=16 | ||
| ✓ | ✓ | ✓ | 76.00 | 63.40 | 2622.32 | 632.94 | 4141.67 | |
[…]
| ✓ | ✓ | ✓ | 61.61 | 52.04 | 2678.45 | 782.86 | 4102.77 | |
[…]
| ✓ | ✓ | ✓ | ✓ | 53.90 | 45.36 | 2579.96 | 668.60 | 4076.78 |
To thoroughly evaluate the impact of our design choices, we conducted several ablation studies. The results are summarized in Tab. 2.
Semantic and visual stream.
First, we examined the individual contributions of the semantic and visual streams to the diffusion process. We trained two model variants, each utilizing only one stream to inject additional context. Despite both variants being provided with an extended context, neither improved upon the baseline results. This limited improvement likely stems from DynamiCrafter being originally trained under matching conditions. In contrast, combining both semantic and visual streams significantly enhanced performance, highlighting their complementary interaction.
3D awareness.
Next, we evaluated the effectiveness of our method’s 3D awareness, achieved through the epipolar cross-attention mechanism. Replacing epipolar cross-attention with standard (vanilla) cross-attention, allowing unrestricted feature aggregation from all tokens, still yielded a considerable improvement of FVD points over the baseline. This model variant, still, demonstrates a significant improvement on the baseline by points in the FVD score but fails to match the performance of the 3D-aware model variant. This can be attributed to the model still leveraging the additional context for the generation but failing to reject features from invalid positions, as seen in Fig. 6, especially when context frames provide minimal additional information due to them being sampled from distant regions.
Temporal awareness.
Further, we assess the effect of temporal embeddings integrated into semantic and visual streams. Removing temporal embeddings results in a performance decline, although still outperforming CamI2V considerably. The temporal embeddings, particularly within the visual stream, explicitly guide the temporal attention of the U-Net to properly interpret timestep-specific context. Without this guidance, the epipolar cross-attention timestep-wise embedded context may be interpreted freely, resulting in impaired performance.
| Sample Range | FVD (VideoGPT) | MSE |
| 45.63 | 2579.96 | |
| 44.21 | 2474.28 | |
| Furthest | 48.52 | 2668.91 |
Context sampling.
Lastly, Tab. 3 compares different sampling strategies for additional context views. Our default method samples context frames from the interval following the generated video. Furthermore, we investigate two extremes: first, sampling a completely unrelated frame, the furthest frame, as shown in Fig. 6. Our results show that this only slightly degrades the visual quality, indicating that our method effectively rejects unrelated features through the induced 3D awareness of the epipolar cross-attention. Second, sampling a frame directly following the video, providing a maximal amount of information to the diffusion process. This only slightly improves our method, showing that it can effectively gather context from loosely placed context views. The qualitative results in Fig. 6 show that our context-aware encoder effectively sorts out unrelated information and provides the diffusion process only with the necessary context.
5.5 Comparison to Novel View Synthesis Approach
We finally compare our method against FrugalNeRF [9], a state-of-the-art novel view synthesis approach for sparse views. Such NeRF-based methods typically rely on an initial sparse 3D reconstruction step, limiting their applicability to diverse scenes. In a two-view setup, the COLMAP pipeline [16], typically employed in this step, only achieves a reliable registration in over of the cases if the frames distance is between and , heavily limiting such approaches if the context frames are too close or too distant.
We train FrugalNeRF on test scenes from the RealEstate10K dataset using the first and th frame as training frames and the intermediate ones as test frames, similar to the End+1 evaluation setup of the context sampling study in Sec. 5.4. Accordingly, we define the reference and context frame for our approach.
Tab. 4 shows that our method slightly outperforms FrugalNeRF while being more broadly applicable, as it does not require a prior reconstruction step. Moreover, because NeRF-based methods typically employ a test-time optimization training scheme, the time needed to learn and render the 3D representation () is substantially larger than the pure rendering time for our approach (), which can be directly applied to novel scenes after training. See the supplementary material for additional details.
| Method | SSIM | LPIPS | PSNR | Time (s) |
| FrugalNeRF | 0.737 | 0.156 | 23.76 | 1265.87 |
| Ours | 0.741 | 0.128 | 24.32 | 8.01 |
6 Conclusion
This paper introduces CamC2V, a novel conditioning mechanism that provides the diffusion process with extensive contextual information derived from multiple context views. Unlike conventional image-to-video diffusion models, which typically rely on a single reference image, our proposed method employs a Context-aware Encoder that encodes additional context through a high-level semantic stream and a 3D-aware visual stream, generating a global semantic representation and a dense, pixel-wise visual embedding from context views. This encoder effectively aggregates relevant features complementing the generation while filtering out unrelated information. As a result, our approach significantly improves visual quality and enhances the accuracy of the generated camera trajectories compared to existing baseline methods.
Supplementary Material
7 Computational Demand
Table 5 reports peak GPU memory and end-to-end inference latency for the baseline models and for our method. As shown, the Context-aware Encoder introduces only a small computational overhead relative to the base model. The visual stream exhibits a larger (yet still minor) latency than the semantic stream because it employs additional timestep- and pixel-wise query tokens to produce a dense visual representation.
| Method | # Params | GPU Memory (GiB) | Latency (s) | |
| Trainable | Total | |||
| DynamiCrafter | - | 2.6 B | 10.43 | 5.143 |
| MotionCtrl | - | 2.6 B | 10.49 | 4.849 |
| CameraCtrl | - | 2.8 B | 11.28 | 4.898 |
| CamI2V | - | 2.9 B | 11.21 | 7.976 |
| CamC2V | 97.4 M | 2.9 B | 11.68 | 8.208 |
| – Semantic Stream | 50.9 M | 50.9 M | 0.19 | 0.004 |
| – Visual Stream | 46.5 M | 46.5 M | 0.17 | 0.043 |
8 Camera Evaluation
We assess the quality of the generated camera trajectories by estimating camera poses with the GLOMAP pipeline. The non-default configuration used in our evaluation is listed in Table 7.
| Parameter | Value |
| ImageReader.single_camera | 1 |
| ImageReader.camera_model | SIMPLE_PINHOLE |
| ImageReader.camera_params | |
| SiftExtraction.estimate_affine_shape | 1 |
| SiftExtraction.domain_size_pooling | 1 |
| SiftMatching.guided_matching | 1 |
| SiftMatching.max_num_matches | 65536 |
| RelPoseEstimation.max_epipolar_error | 4 |
| BundleAdjustment.optimize_intrinsics | 0 |
9 Additional Qualitative Results
To further assess the generative quality we provide further qualitative results of our method in Fig. 9 compared to the baseline method CamI2V.
10 COLMAP Evaluation
NeRF or 3D Gaussian Splatting-based approaches often rely on an initial sparse 3D reconstruction step to get an initial pointcloud of the scene. Typically, COLMAP [16] is used as a sparse 3D reconstruction pipeline. The pipeline can be split in three stages, namely feature extraction, feature matching and point triangulation. After the first stage, we insert the ground-truth camera poses into the pipeline to improve the triangulation performance and kepp it fair to our model, which is also provided with ground-truth camera poses.
| Parameter | Value |
| ImageReader.single_camera | 1 |
| ImageReader.camera_model | SIMPLE_PINHOLE |
| ImageReader.camera_params | |
| Mapper.tri_ignore_two_view_tracks | 0 |
| Mapper.num_threads | 16 |
| Mapper.init_min_tri_angle | 1 |
| Mapper.multiple_models | 0 |
| Mapper.extract_colors | 0 |
11 Novel View Synthesis Qualitative Results
We provide additional qualitative results comparing our model against FrugalNeRF in Fig. 9. It is visible that FrugalNeRF not only suffers from an insufficient initial sparse reconstruction step, indicated by the black borders, but also introduces artifacts, especially for fine-grained details and edges, where the 3D representation is not sufficiently learned. Overall, our approach mitigates such artifacts and produces clearer more realistic looking images.
In Fig. 8, we show the success rate of this pipeline plotted against the distance between the two used frames. All failure cases come from the triangulation stage, where the model fails to confidently match and triangulate points between images. Especially, for larger frame distances the success rate significantly drops, eventhough the images show sufficient overlap in our visual inspection. This ultimately only leaves a small window from to where the pipeline is able to reliably produce a sparse 3D reconstruction in over of the cases. Overall, in our used evaluation setup using the first and th frame the pipeline has a success rate of .